OpenStudy (samirahdanyel):

Using technology, determine the monthly payment on a 6 year loan of $15,250 at 3.5% compounded monthly. Round your answer to the nearest cent. a. $234.75 c. $582.70 b. $235.13 d. $590.05

OpenStudy (mathmate):

It is not necessary to use technology. Using a calculator, you can calculate the following: Monthly Payment Amount =P*R^n*(i/12)/(R^n-1) where P is principal, or amount borrowed i=annual interest (in decimal, not percent) R=1+i/12=gross amount with interest after one period (month). It comes up to one of the smaller amounts. If you use technology, it should be even simpler, but you have to make sure you use the tool according to instructions.